# Quantumness certification via non-demolition measurements

### v1

@@ IV.4 Connection and comparison with Leggett-Garg inequalities
In this section, we compare the procedure for detecting MR violations via the QNDM protocol with the usual tool employed so far to identify the violation of MRps, i.e., the Leggett-Garg inequalities. The LGIs are usually formulated for binary observables [5, 8, 10, 6], namely for observables that can have two outcomes. Hence, we now assume that the eigenvalues of the observable $\hat{A}$ are $a_{i}=\pm 1$ for any time $t_{i}$. LGIs are built using quantum correlators at two times, say $t_{i}$ and $t_{j}$, which are defined as [8, 10, 6]
$C_{ij}=\frac{1}{2}{\rm Tr}\Big[\big(\hat{A}^{H}(t_{i})\hat{A}^{H}(t_{j})+\hat{A}^{H}(t_{j})\hat{A}^{H}(t_{i})\big)\hat{\rho}^{0}\Big]$ (30)
with $\hat{A}^{H}(t_{i})=\hat{U}^{\dagger}(t_{i},0)\hat{A}\hat{U}(t_{i},0)$ represented in Heisenberg picture. The expression of $C_{ij}$ can be rewritten as
$C_{ij}=\sum_{i,j,k}a_{i}a_{j}P(a_{k},a_{j},a_{i})=\sum_{i,j,k}a_{i}a_{j}P^{(0)}_{i}P^{(1)}_{i\rightarrow j}P^{(2)}_{j\rightarrow k}=\sum_{i,j,k}a_{i}a_{j}P_{\rm QNDM}(k,j,j,i,i),$ (31)
where $P(a_{k},a_{j},a_{i})=P^{(0)}_{i}P^{(1)}_{i\rightarrow j}P^{(2)}_{j\rightarrow k}=P_{\rm QNDM}(k,j,j,i,i)$ is the joint probability to record the measurement outcomes $a_{k},a_{j},a_{i}$ at three consecutive times using a sequential scheme of projective measurements. Following Ref. [10], we introduce the Leggett-Garg (LG) parameter
@@
that reads as $K=1-4\left[P(1,-1,1)+P(-1,1,-1)\right]$ and obeys the LGI $3\leq K\leq 1$ if the system under consideration is classical and fulfils all MR assumptions. This means that, if the $3\leq K\leq 1$ is not satisfied, then at least one condition of MR is violated. However, the opposite is not guaranteed because there are cases where the MR is violated, but the value of $K$ is still included in the range $3\leq K\leq 1$.
Starting from the QNDM quasi-probability density function (22), the correlators $C_{ij}$ can be equivalently expressed as [6, 26]
$C_{ij}=\sum_{i,j}a_{i}a_{j}P_{\rm QNDM}(j,i),$ (33)
where
$P_{\rm QNDM}(j,i)$ $=$ $\sum_{l,m,k}\left[P_{\rm QNDM}(k,j,m,i,l)+P_{\rm QNDM}(k,m,j,i,l)+\right$. (34)
$+$ $\left.P_{\rm QNDM}(k,j,m,l,i)+P_{\rm QNDM}(k,m,j,l,i)\right]$
are Kirkwood-Dirac quasi-probabilities [48, 25, 63], as commented in Sec. IV.2.1. From Eq. (33), we can evince that the correlators entering the LGI $3\leq K\leq 1$ are functions of QNDM quasiprobabilities at three times, whose distribution thus has full information about the LG parameter $K$. Accordingly, using the QNDM quasi-probability density function, one can directly single out the value of $K$ and understand whether the LGI is violated.
We now report the expression of the LG parameter $K$ as a function of the QNDM quasi-probabilities at three times; the derivation of this result is in Ref. [26]. Since we are working with binary variables, each of them can take two values: $a_{i}$ and $a_{i}$. For simplicity, we refer to them with the indices $i$ and $\bar{i}$; for example, $P_{\rm QNDM}(\bar{k},j,m,i,\bar{i})$ stands for $P_{\rm QNDM}(-a_{k},a_{j},a_{m},a_{i},-a_{i})$. Using Eq. (33), the LG parameter $K$ of Eq. (32) reads as
@@
with $f(k,j,m,i,l)=(a_{i}+a_{l})(a_{j}+a_{m})+a_{k}(a_{j}+a_{m}-a_{i}-a_{l})$. It is convenient to separate $K$ in its classical and quantum contributions: $K=K_{cl}+K_{q,1}+K_{q,2}$. The classical contribution is obtained when $(m=j,l=i)$, while the quantum ones $K_{q,1}$ and $K_{q,2}$ correspond to $(m=j,l=\bar{i}\neq i)$ and $(m=\bar{j}\neq j,l=i)$, respectively. It can be shown [26] that $K_{cl}$ satisfies the LGI, i.e., $3\leq K_{cl}\leq 1$, and that $K_{q,1}=0$. Therefore, when present, the violation of the LGI is due to the second quantum contribution $K_{q,2}$ of $K$, which is equal to
@@
From this analysis, we can learn that the contributions with $m=\bar{j}\neq j$ and $l=i$, entailed by the superposition between eigenstates of $\hat{A}$, are the ones responsible for the violation of the LGI. This implies, in turn, that whenever the MRps is violated but the LGI is satisfied, we are dealing with a combination of initial state, system dynamics, and measurement observable such that $K_{q,2}$ vanishes. In such a case, the total quantum contribution $K_{q}$ vanishes as well, $K_{q}=0$, with the result that $K=K_{cl}$, confirming that the violation of the LGI is only a sufficient condition for identifying MR violations. Indeed, there could be trajectories in which $l\neq i$, meaning that the system is initially in a superposition of $\hat{A}$’s eigenstates that naturally violates the MRps. These trajectories cannot be identified by LGI since $K_{q,1}=0$. On the contrary, if we use $\mathcal{P}_{\rm QNDM}(\Delta)$, the terms $P_{\rm QNDM}(k,j,j,i,\bar{i})$ in (22) always contribute to negative regions of the distribution, whose presence is both a sufficient and necessary condition for the violation of the MRps and thus of MR.
Overall, detecting violations of MR using the framework of NSIT conditions provides a significantly robust criterion [27, 28]. However, for complex quantum systems with many degrees of freedom or strongly correlated dynamics, verifying whether the NSIT conditions are fulfilled can be very challenging. Many solutions proposed so far in the literature [27, 28, 29, 30] to evaluate MR do not work in general, except for a specific choice of parameters, initial state, and measurement times. These constraints are not present in the QNDM approach, which can therefore represent an operational protocol for quantumness certification, at the price, however, of using an auxiliary system acting as a detector.
Figure 2: Comparison between the prediction of the LGI and QNDM approach. In the top-left panel, we plot the LG parameter $K$ as a function of $\omega\tau$. The shaded regions are the ones for which the LGI is violated. The other panels show the quasi-probability distribution $\mathcal{P}_{\rm QNDM}(\Delta)$ evaluated at three different values of $\omega\tau$. The distribution $\mathcal{P}_{\rm QNDM}(\Delta)$ always shows negative regions, while only for the green circle point the LGI is violated.
@@ IV.4.1 Example
In order to clarify how the QNDM approach works and performs against the traditional LGIs, we report a specific example (present also in Refs. [6, 26]). We consider a two-level quantum system that is initialized in the state $\mathinner{|{\psi_{0}}\rangle}=(\mathinner{|{0}\rangle}+i\mathinner{|{1}\rangle})/\sqrt{2}$ where $\mathinner{|{0}\rangle}$ and $\mathinner{|{1}\rangle}$ are the eigenstates of the Pauli operator $\hat{\sigma}_{z}$. The dynamics is generated by the Hamiltonian $\hat{H}=\omega\hat{\sigma}_{x}/2$, and we measure the operator $\hat{A}=\hat{\sigma}_{z}$ at times $t_{0}=0$, $t_{1}=\tau$ and $t_{2}=2\tau$.
The correlators $C_{ij}$ and the corresponding LG parameter $K$ can be calculated analytically, such that $K=2\cos(\omega\tau)-\cos(2\omega\tau)$. As discussed above, any value of $K$ outside the range [3,1][-3,1] implies the violation of at least one MR condition; instead, it is not guaranteed that the MR holds if $3\leq K\leq 1$.
The numerical results for the example are shown in Fig. 2. The top-left panel shows the value of $K$ as a function of $\omega\tau$ with the shaded regions representing the situations in which the LGI is violated. The colored shapes (dot, square, triangle) give the value of $K$ for three specific values of $\omega\tau$. The other panels of Fig. 2 depict the QNDM quasi-probability distribution associated with each colored shape. In correspondence of the green circle, the LGI is violated and the corresponding QNDM distribution $\mathcal{P}_{\rm QNDM}(\Delta)$ has negative regions. However, for the other two colored shapes (the orange square and the purple circle), a more interesting fact occurs: while the LGI is satisfied, $\mathcal{P}_{\rm QNDM}(\Delta)$ presents negative regions, and it allows us to identify quantum features of the system even when the LGI fails. From the analytical expression of $K$, the violation of the LGI occurs for $0\leq\omega\tau\leq\pi/2$ and $3\pi/2\leq\omega\tau\leq 2\pi$. Hence, in this example, the LGI correctly identifies the violation of MRps in only half of the cases, while the QNDM protocol always succeeds.
@@ V Resources for realizing QNDM, and robustness against environmental noise
As discussed in Secs. IV.3 and IV.4, the QNDM approach allows for the identification of the violation of MR even when the LGI fails. It is thus natural to ask how many resources one needs to realize a sequence of non-demolition measurements, and compare them to the resources required to reconstruct the LGI. Here, we refer to the protocols discussed in Sec. IV.4, including the example in IV.4.1, by performing a detailed comparison of the implementation costs following Ref. [94].
@@
Figure 3: a) Simulation results for the computation of the LG parameter $K=C_{01}+C_{12}-C_{02}$ in a scenario where the measurement statistical uncertainty is present, with $N_{\text{shots}}=10^{4}$. The solid blue line represents the simulated average curve (obtained over $N_{\text{shots}}$ repetitions), while the dashed red line corresponds to the theoretical prediction as given in the example at the end of Sec. IV.4. The blue-shaded region origins from the statistical uncertainty, while the yellow-shaded regions highlight the range of $\omega\tau$ where, considering statistical errors, a violation of the LGI is confidently detected. b) Plot of the quasi-probability density functions $\mathcal{P}_{\rm QNDM}(\Delta)$ for $\delta\lambda=1$ and $N_{\text{shots}}=10^{2}$. These curves are obtained by taking the inverse Fourier transform of the simulated data for $\omega\tau=1.5$, and the orange error bars represent the statistical uncertainty. Notice that, in order to extract single probability values from $\mathcal{P}_{\rm QNDM}(\Delta)$, we must integrate it over a fixed interval of $\Delta$. The negative region in $\mathcal{P}_{\rm QNDM}(\Delta)$ is visible for any value of $\omega\tau$. These panels are adapted from figures in Ref. [94].
@@
Now, let us focus on how statistical uncertainty, due to a finite number of experimental repetitions, affects the accuracy with which a MR violation can be identified for both the LG and QNDM protocols. In Fig. 3 panel a), we show the LG parameter $K$ obtained by computing the correlators $C_{01},C_{12},C_{02}$ in simulation with $N_{\text{shots}}=10^{4}$, while in panel b) we plot the QNDM quasi-probability density function $\mathcal{P}_{\rm QNDM}(\Delta)$, using $N_{\text{shots}}=10^{2}$ experimental repetitions with $\delta\lambda=1$ (thus, $N_{\lambda}=10^{2}$). This means that, to run the experiments, we need similar resources: $N_{\text{LG}}=3\cdot 10^{4}$ and $N_{\text{QNDM}}=2\cdot 10^{4}$.
@@
In panel a) of Fig. 3, the blue shadowed regions are delineated by the variance associated with the statistical uncertainty of measurements, while the yellow-shaded regions represent the interval of $\omega\tau$ where we are “statistically certain” to violate the LGI. The dashed red curve represents the ideal value of $K$ (i.e., with no statistical uncertainty), while the solid blue curve is the mean value of $K$ averaged over the experimental repetitions. In the ideal case, violations of the LGI occur in the ranges $0\leq\omega\tau\leq\pi/2$ and $3\pi/2\leq\omega\tau\leq 2\pi$, covering $50\%$ of the total $\omega\tau$ range. The presence of measurement statistical uncertainty significantly reduces the confidence interval for the LGI’s violation to $38\%$ of the total $\omega\tau$ range [94].
@@
Panel b) of Fig. 3 shows the QNDM quasi-probability density function obtained with $\omega\tau=1.5$. According to the QNDM protocol, we need to take into account many values of $\mathcal{G}_{\lambda}$; in this case, we take $0\leq\lambda\leq 100$ with discretization $\delta\lambda=1$. As we can see from the figure, the negative region of $\mathcal{P}_{\rm QNDM}(\Delta)$ around $\Delta=2$ is clearly visible, identifying the presence of quantum effects. It turns out that the negative region is visible for any value of $\omega\tau$, albeit in the presence of a measurement statistical uncertainty. With these parameters, the QNDM protocol identifies the violation of MR for any $\omega\tau$, requiring a comparable amount of resources to those needed to compute the LG parameter $K$. We conclude that, with the same amount of resources, the QNDM always identifies the violation of MR while the LGI fails in about $60\%$ of the cases.
@@
Panel a) of Fig. 4 refers to $N_{\text{shots}}=10^{4}$ repetitions, for a total of $N_{\text{LG}}=3\cdot 10^{4}$ repetitions. From the figure, we can see that the regions where the LGI is violated are further shrunk to $24\%$ of the total $\omega\tau$ range, respectively. These results show that the use of the LGI as a witness of an MR violation may be problematic in the presence of environmental noise and gate/control errors.
Panel b) of Fig. 4 displays the QNDM quasi-probability density function obtained with $N_{\text{shots}}=10^{2}$ and $\delta\lambda=1$ for a total of $N_{\text{QNDM}}=2\cdot 10^{4}$ repetitions. Despite a quite low level of resources and the presence of environmental noise, the negative region of $\mathcal{P}_{\rm QNDM}(\Delta)$ is still clearly visible. Therefore, in this case, the advantage of using QNDM is even more pronounced since, with a fully comparable amount of resources, they allow for identifying the violation of the MR, while the LGI fails roughly in $75\%$ of cases.

### v2

@@ IV.4 Connection and comparison with Leggett-Garg inequalities
In this section, we compare the QNDM protocol with the usual tool employed so far to identify the violation of MRps, i.e., the Leggett-Garg inequalities. The LGIs are usually formulated for binary observables [5, 85, 83, 6], namely for observables that can have two outcomes. Hence, we now assume that the eigenvalues of the observable $\hat{A}$ are $a(t_{n})=\pm 1$ for any time $t_{n}$. LGIs are built using quantum correlators at two times, say $t_{n}$ and $t_{m}$, which are defined as [85, 83, 6]
$C_{nm}=\frac{1}{2}{\rm Tr}\Big[\big(\hat{A}^{H}(t_{n})\hat{A}^{H}(t_{m})+\hat{A}^{H}(t_{m})\hat{A}^{H}(t_{n})\big)\hat{\rho}^{0}\Big]$ (31)
with $\hat{A}^{H}(t)=\hat{U}^{\dagger}(t,0)\hat{A}\hat{U}(t,0)$ represented in Heisenberg picture. The expressions of $C_{nm}$ with $n,m=0,1,2$ and $n<m$ are respectively equal to
$C_{01}$ $=$ $\sum_{i,j,k}a_{i}a_{j}P(a_{k},a_{j},a_{i})$ (32)
$C_{12}$ $=$ $\sum_{i,j,k}a_{j}a_{k}P(a_{k},a_{j},a_{i})$ (33)
$C_{02}$ $=$ $\sum_{i,j,k}a_{i}a_{k}P(a_{k},a_{j},a_{i}),$ (34)
where $P(a_{k},a_{j},a_{i})=P^{(0)}_{i}P^{(1)}_{i\rightarrow j}P^{(2)}_{j\rightarrow k}=P_{\rm QNDM}(k,j,j,i,i)$ is the joint probability to record the measurement outcomes $a_{i},a_{j},a_{k}$ at times $t_{0},t_{1},t_{2}$ using a sequential scheme of projective measurements. Following Ref. [83], we introduce the Leggett-Garg (LG) parameter
@@
that reads as $K=1-4\left[P(1,-1,1)+P(-1,1,-1)\right]$ and obeys the LGI $3\leq K\leq 1$ if the system under consideration is classical and fulfils all the assumptions of MR. This means that, if the $3\leq K\leq 1$ is not satisfied, then at least one condition of MR is violated. However, the opposite is not guaranteed because there are cases where the MR is violated, but the value of $K$ is still included in the range $3\leq K\leq 1$.
Starting from the QNDM quasi-probability density function (22), for binary observables the correlators $C_{nm}$ can be equivalently expressed as [6, 26]
$C_{nm}=\sum_{u,v}a_{u_{n}}a_{v_{m}}P_{\rm QNDM}(v_{m},u_{n}),$ (36)
where $a_{u_{n}}\mathrel{\mathop{\ordinarycolon}}=a_{u}(t_{n})$, $a_{v_{m}}\mathrel{\mathop{\ordinarycolon}}=a_{v}(t_{m})$, and $P_{\rm QNDM}(v_{m},u_{n})$ are Kirkwood-Dirac quasi-probabilities [54, 25, 68], as commented in Sec. IV.1.1. For example, $C_{01}=\sum_{i,j}a_{i}a_{j}P_{\rm QNDM}(j,i)$ with
$P_{\rm QNDM}(j,i)=\sum_{l,m,k}\left[P_{\rm QNDM}(k,j,m,i,l)+P_{\rm QNDM}(k,m,j,i,l)+P_{\rm QNDM}(k,j,m,l,i)+P_{\rm QNDM}(k,m,j,l,i)\right].$ (37)
From Eq. (36), we thus evince that the LG parameter $K$ can be written as a function of the QNDM quasi-probabilities at three times, whose distribution implicitly contains the information of the two-times correlators $C_{nm}$ and of the three-time correlator $C_{ijk}=\sum_{i,j,k}a_{i}a_{j}a_{k}P(a_{k},a_{j},a_{i})$. Accordingly, using the QNDM quasi-probability density function, one can directly single out the value of $K$ and understand whether the LGI is violated.
We now report and comment on the expression of the LG parameter $K$ as a function of the QNDM quasi-probabilities at three times; the derivation of this result is in Ref. [26]. Since we are working with binary variables, each of them can take two values: $a_{n}$ and $a_{n}$. For simplicity, we refer to them with the indices $n$ and $\bar{n}$; for example, $P_{\rm QNDM}(\bar{k},j,m,i,\bar{i})$ stands for $P_{\rm QNDM}(-a_{k},a_{j},a_{m},a_{i},-a_{i})$. Using Eq. (36), the LG parameter $K$ of Eq. (35) reads as
@@
with $f(k,j,m,i,l)=(a_{i}+a_{l})(a_{j}+a_{m})+a_{k}(a_{j}+a_{m}-a_{i}-a_{l})$. It is convenient to separate $K$ in its classical and quantum contributions: $K=K_{cl}+K_{q,1}+K_{q,2}$. The classical contribution $K_{cl}$ is obtained when $(m=j,l=i)$, while the quantum ones $K_{q,1}$ and $K_{q,2}$ correspond to $(m=j,l=\bar{i}\neq i)$ and $(m=\bar{j}\neq j,l=i)$, respectively. It can be shown [26] that $K_{cl}$ satisfies the LGI, i.e., $3\leq K_{cl}\leq 1$, and that $K_{q,1}=0$. Therefore, when present, the violation of the LGI is due to the second quantum contribution $K_{q,2}$ of $K$, which is equal to
@@
The contributions with $m=\bar{j}\neq j$ and $l=i$ entering $K_{q,2}$ are entailed by some superpositions between eigenstates of $\hat{A}$, thus violating MRps. This implies, in turn, that whenever the MRps is violated but the LGI is satisfied, we are dealing with a combination of initial state, system dynamics, and measurement observable such that $K_{q,2}$ vanishes. In such a case, the total quantum contribution $K_{q}$ vanishes as well, $K_{q}=0$, with the result that $K=K_{cl}$, confirming that the violation of the LGI is only a sufficient condition for identifying violations of MRps. Indeed, there could be trajectories in which $l\neq i$ (thus with a superposition of $\hat{A}$’s eigenstates) that are not identified by LGI since $K_{q,1}=0$. On the contrary, if we take into account $\mathcal{P}_{\rm QNDM}(\Delta)$, the terms $P_{\rm QNDM}(k,j,j,i,\bar{i})$ in (22) always contribute to negative regions of the distribution, whose presence is a sufficient and necessary condition for the violation of MRps.
Overall, the QNDM approach does not have to face with constraints on the choice of parameters, initial state, and measurement times, as it occurs for standard methods in the literature [86, 87, 88, 84] that address the problem of detecting the presence of quantum superposition states, in the initial system’s state and/or generated dynamically. From an operational standpoint, QNDM can thus represent an efficient and robust protocol for certifying quantumness, albeit at the cost of using an auxiliary system that acts as a detector.
Figure 2: Comparison between the prediction of the LGI and QNDM approach. In the top-left panel, we plot the LG parameter $K$ as a function of $\omega\tau$. The shaded regions are the ones for which the LGI is not satisfied. The other panels show the quasi-probability distribution $\mathcal{P}_{\rm QNDM}(\Delta)$ evaluated at three different values of $\omega\tau$. The distribution $\mathcal{P}_{\rm QNDM}(\Delta)$ always shows negative regions, while only for the green circle point the LGI is violated.
@@ IV.4.1 Example
Let us consider a two-level quantum system that is initialized in the state $\mathinner{|{\psi_{0}}\rangle}=(\mathinner{|{0}\rangle}+i\mathinner{|{1}\rangle})/\sqrt{2}$ where $\mathinner{|{0}\rangle}$ and $\mathinner{|{1}\rangle}$ are the eigenstates of the Pauli operator $\hat{\sigma}_{z}$. The dynamics is generated by the Hamiltonian $\hat{H}=\omega\hat{\sigma}_{x}/2$, and we measure the operator $\hat{A}=\hat{\sigma}_{z}$ at times $t_{0}=0$, $t_{1}=\tau$ and $t_{2}=2\tau$ [6, 26].
The two-times correlators $C_{nm}$ and the corresponding LG parameter $K$ can be calculated analytically, such that $K=2\cos(\omega\tau)-\cos(2\omega\tau)$. As discussed above, any value of $K$ outside the range [3,1][-3,1] implies the violation of at least one MR condition; instead, it is not guaranteed that the MR holds if $3\leq K\leq 1$.
The numerical results for the example are shown in Fig. 2. The top-left panel shows the value of $K$ as a function of $\omega\tau$ with the shaded regions representing the situations in which the LGI is not satisfied. The colored shapes (dot, square, triangle) give the value of $K$ for three specific values of $\omega\tau$. The other panels of Fig. 2 depict the QNDM quasi-probability distribution associated with each colored shape. In correspondence with the green circle, the LGI is violated, and the corresponding QNDM distribution $\mathcal{P}_{\rm QNDM}(\Delta)$ has negative regions. However, for the other two colored shapes (the orange square and the purple circle), a more interesting fact occurs: while the LGI is satisfied, $\mathcal{P}_{\rm QNDM}(\Delta)$ presents negative regions, and it allows us to identify quantum features of the system even when the LGI fails. From the analytical expression of $K$, the violation of the LGI occurs for $0\leq\omega\tau\leq\pi/2$ and $3\pi/2\leq\omega\tau\leq 2\pi$. Hence, in this example, the LGI correctly identifies the violation of MRps in only half of the cases, while the QNDM protocol always succeeds.
@@ V Resources for realizing QNDM, and robustness against environmental noise
As discussed in Secs. IV.2 and IV.4, the QNDM approach allows for identifying the presence of quantum features in a system’s state even when the LGI fails. It is thus natural to ask how many resources one needs to realize a sequence of non-demolition measurements, and compare them to the resources required to reconstruct the LGI. Here, we refer to the protocols discussed in Sec. IV.4, including the example in IV.4.1, by performing a detailed comparison of the implementation costs following Ref. [89].
@@
Figure 3: a) Simulation results for the computation of the LG parameter $K=C_{01}+C_{12}-C_{02}$ in a scenario where the measurement statistical uncertainty is present, with $N_{\text{shots}}=10^{4}$. The solid blue line represents the simulated average curve (obtained over $N_{\text{shots}}$ repetitions), while the dashed red line corresponds to the theoretical prediction as given in the example at the end of Sec. IV.4. The blue-shaded region originates from the statistical uncertainty, while the yellow-shaded regions highlight the range of $\omega\tau$ where, considering statistical errors, a violation of the LGI is confidently detected. b) Plot of the quasi-probability density functions $\mathcal{P}_{\rm QNDM}(\Delta)$ for $\delta\lambda=1$ and $N_{\text{shots}}=10^{2}$. These curves are obtained by taking the inverse Fourier transform of the simulated data for $\omega\tau=1.5$, and the orange error bars represent the statistical uncertainty. Notice that, in order to extract single probability values from $\mathcal{P}_{\rm QNDM}(\Delta)$, we must integrate it over a fixed interval of $\Delta$. The negative region in $\mathcal{P}_{\rm QNDM}(\Delta)$ is visible for any value of $\omega\tau$. These panels are adapted from figures in Ref. [89].
@@
Now, let us focus on how statistical uncertainty, due to a finite number of experimental repetitions, affects the accuracy with which a violation of MRps can be identified for both the LG and QNDM protocols. In Fig. 3 panel a), we show the LG parameter $K$ obtained by computing the correlators $C_{01},C_{12},C_{02}$ in simulation with $N_{\text{shots}}=10^{4}$, while in panel b) we plot the QNDM quasi-probability density function $\mathcal{P}_{\rm QNDM}(\Delta)$, using $N_{\text{shots}}=10^{2}$ experimental repetitions with $\delta\lambda=1$ (thus, $N_{\lambda}=10^{2}$). This means that, to run the experiments, we need similar resources: $N_{\text{LG}}=3\cdot 10^{4}$ and $N_{\text{QNDM}}=2\cdot 10^{4}$.
@@
In panel a) of Fig. 3, the blue shadowed regions are delineated by the variance associated with the statistical uncertainty of measurements, while the yellow-shaded regions represent the interval of $\omega\tau$ where we are “statistically certain” not to satisfy the LGI. The dashed red curve represents the ideal value of $K$ (i.e., with no statistical uncertainty), while the solid blue curve is the mean value of $K$ averaged over the experimental repetitions. In the ideal case, violations of the LGI occur in the ranges $0\leq\omega\tau\leq\pi/2$ and $3\pi/2\leq\omega\tau\leq 2\pi$, covering $50\%$ of the total $\omega\tau$ range. The presence of measurement statistical uncertainty significantly reduces the confidence interval for the LGI’s violation to $38\%$ of the total $\omega\tau$ range [89].
@@
Panel b) of Fig. 3 shows the QNDM quasi-probability density function obtained with $\omega\tau=1.5$. According to the QNDM protocol, we need to take into account many values of $\mathcal{G}_{\lambda}$; in this case, we take $0\leq\lambda\leq 100$ with discretization $\delta\lambda=1$. As we can see from the figure, the negative region of $\mathcal{P}_{\rm QNDM}(\Delta)$ around $\Delta=2$ is clearly visible, identifying the presence of quantum effects. It turns out that the negative region is visible for any value of $\omega\tau$, albeit in the presence of a measurement statistical uncertainty. With these parameters, the QNDM protocol certifies the violation of MRps for any $\omega\tau$, requiring a comparable amount of resources to those needed to compute the LG parameter $K$. We conclude that, with the same amount of resources, the QNDM always succeeds while the LGI fails in about $60\%$ of the cases.
Panel a) of Fig. 4 refers to $N_{\text{shots}}=10^{4}$ repetitions, for a total of $N_{\text{LG}}=3\cdot 10^{4}$ repetitions. From the figure, we can see that the regions where the LGI is not satisfied are further shrunk to $24\%$ of the total $\omega\tau$ range, respectively. On the other hand, panel b) of Fig. 4 displays the QNDM quasi-probability density function obtained with $N_{\text{shots}}=10^{2}$ and $\delta\lambda=1$ for a total of $N_{\text{QNDM}}=2\cdot 10^{4}$ repetitions. Despite a quite low level of resources and the presence of environmental noise, the negative region of $\mathcal{P}_{\rm QNDM}(\Delta)$ is still clearly visible. Therefore, in this case, the advantage of using QNDM is even more pronounced since, with a fully comparable amount of resources, they always allow for quantumness certification, while the LGI fails roughly in $75\%$ of cases.